a new car depreciates by 20% in its first year. Each succeeding year it depreciates by 10%. Assume the car is still in good working condition after 10 years. If the car costs $25000, calculate te value of the car after 10 years. Express your answer to the nearest $10
Answer by mathstutor458(57)

the car costs $25000
car depreciates by 20% in its first year=0.2
means car value after first year = 25000 * (1-0.2)= 25000 * 0.8
Each succeeding year it depreciates by 10% = 10/100 =0.1
car value after second year = 25000 * 0.8 * (1-0.1)= 25000 * 0.8* 0.9
Each succeeding year it depreciates by 10% ,
So,
car value after 10 years = 25000 * 0.8 * (0.9)^9 (one year = 0.8,all the nine years=0.9, so i write (0.9)^9)
=$7,748.41------>answer